OpenStudy (anonymous):

Estimate 1 4 8-- x 2-- 3 5 using any method you like. Then find the actual product. Use your estimate to show that your answer is reasonable.

OpenStudy (lgbasallote):

\[\huge 8 \frac 13 \times 2 \frac 45?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):